Solve the following for x: (e^4x) - (2e^2x) + 1=0
Answer by lwsshak3(11628)

Solve the following for x:
(e^4x) - (2e^2x) + 1=0
let u=e^2x
u^2=e^4x
..
u^2-2u+1=0
(u-1)^2=0
u=1=e^2x
e^2x=1
take log of both sides
2xlne=ln1
lne=1
ln1=0
2x=0
x=0